# Adaptive Smooth Non-Stationary Bandits

### v1

@@ Abstract
Next, we show this optimal dynamic regret can be attained adaptively, without knowledge of $\beta,\lambda$. To contrast, even with parameter knowledge, upper bounds were only previously known for limited regimes $\beta\leq 1$ and $\beta=2$ (Slivkins, 2014; Krishnamurthy and Gopalan, 2021; Manegueu et al., 2021; Jia et al., 2023). Thus, our work resolves open questions raised by these disparate threads of the literature.
We also study the problem of attaining faster gap-dependent regret rates in non-stationary bandits. While such rates are long known to be impossible in general (Garivier and Moulines, 2011), we show that environments admitting a safe arm (Suk and Kpotufe, 2022) allow for much faster rates than the worst-case scaling with $\sqrt{T}$. While previous works in this direction focused on attaining the usual logarithmic regret bounds, as summed over stationary periods, our new gap-dependent rates reveal new optimistic regimes of non-stationarity where even the logarithmic bounds are pessimistic. We show our new gap-dependent rate is tight and that its achievability (i.e., as made possible by a safe arm) has a surprisingly simple and clean characterization within the smooth Hölder class model.
@@ 1   Introduction
In the multi-armed bandit (MAB) problem, an agent sequentially chooses actions, from a set of $K$ arms, based on partial and uncertain feedback in the form of (bounded) rewards $Y_{t}(a)$ for past actions $a\in[K]$ (see Bubeck and Cesa-Bianchi, 2012; Slivkins, 2019; Lattimore and Szepesvári, 2020, for general surveys). The goal is to maximize the cumulative reward.
We consider the non-stationary variant of the problem, where rewards are obliviously adversarial. In particular, we consider a smooth reward model of non-stationary MAB where only mild Hölder class assumptions are made on changes in rewards over time. In fact, this model captures any finite-horizon bandit problem (e.g., via a polynomial interpolation). Additionally, the degree of smoothness (as measured by the Hölder exponent or coefficient of the associated Hölder class) can be considered a more fine-grained measure of non-stationarity in comparison to conventional measures appearing in other works on non-stationary MAB. Indeed, the rates in this model smoothly interpolate between the more parametric $\sqrt{LT}$ rates seen in switching bandits (Garivier and Moulines, 2011), with $L$ switches in rewards over horizon $T$, and the $V^{1/3}T^{2/3}$ rates in terms of the total variation measure $V$ quantifying magnitude of total changes in rewards (Besbes et al., 2019).
@@
Our result is also somewhat surprising since prior approaches (Krishnamurthy and Gopalan, 2021; Manegueu et al., 2021; Jia et al., 2023) more or less relied on confidence intervals on the magnitude of change in rewards, whose design requires knowledge of the smoothness. In non-parametric statistics, it’s long been known that it’s impossible to design confidence intervals adaptive to unknown smoothness (Low, 1997), ruling out approaches of this kind.
@@ 1.1   Further Discussion on Related Works
Switching bandits was first considered in the adversarial setting by Auer et al. (2002), where a version of EXP3 was shown to attain optimal dynamic regret $\sqrt{LT}$ when tuned with knowledge of the number $L$ of switches. Later works showed similar guarantees in this problem for procedures inspired by stochastic bandit algorithms (Kocsis and Szepesvári, 2006; Yu and Mannor, 2009; Garivier and Moulines, 2011; Mellor and Shapiro, 2013; Liu et al., 2018; Cao et al., 2019). Recently, Auer et al. (2018, 2019); Chen et al. (2019) established the first adaptive and optimal dynamic regret guarantees, without requiring knowledge of $L$. Other non-stationarity measures, such as the aforementioned total variation, or more nuanced counts than $L$ were recently studied (Suk and Kpotufe, 2022; Abbasi-Yadkori et al., 2023).
@@ 1.2   Contributions
Works on Smooth BanditsAdaptive?ParametersDynamic Regret Upper BoundManegueu et al. (2021) Noβ(0,1]Tβ+12β+1λ12β+1Kβ2β+1
Slivkins (2014)
Krishnamurthy and Gopalan (2021)
 Noβ=1Tλ13K13
Jia et al. (2023) Noβ=1,2,K=2Tβ+12β+1λ12β+1
This Work
@@
Slivkins (2014)
Krishnamurthy and Gopalan (2021)
@@ 2.1   Preliminaries and Notation
Arm $a$ at round $t$ has random reward $Y_{t}(a)\in[0,1]$ with mean $\mu_{t}(a)$. A (possibly randomized) algorithm $\pi$ selects at each round $t$ some arm $\pi_{t}\in[K]$ and observes reward $Y_{t}(\pi_{t})$. The goal is to minimize the dynamic regret, i.e., the expected regret to the best arm at each round. This is defined as:
@@
We’ll use $R_{\mathcal{E}}(\pi,T)$ to denote the expected regret under an environment $\mathcal{E}$.
@@
Throughout this paper, in theorem statements we’ll use $C_{0},C_{1},\ldots$ to denote universal constants free of $K,T,\beta,\lambda,\{\mu_{t}(a)\}_{t\in[T],a\in[K]}$. In proofs, universal constants $c_{0},c_{1},\ldots$ will be used.
@@ 2.2   Smooth Non-Stationary Bandits
We say a bandit environment is $(\beta,\lambda)$-Hölder if, for every arm $a\in[K]$, there exists a $(\beta,\lambda)$-Hölder function $f$ such that the gap function (in time) is realized by $f$, i.e. $\delta_{t}(a)=f(t/T)$ for all $t\in[T]$. We’ll use $\Sigma(\beta,\lambda)$ to denote the class of bandit environments which are $(\beta,\lambda)$-Hölder over $T$ rounds.
@@ 3   Dynamic Regret Lower Bound
We first characterize the minimax regret rate over the class of problems in $\Sigma(\beta,\lambda)$. For comparison, Jia et al. (2023, Theorem 3.4) already established a lower bound for integer smoothness $\beta\in\mathbb{Z}_{\geq 1}$, $K=2$ arms, and fixed Hölder coefficient $\lambda=1$. Our main novelty here is to show a more comprehensive lower bound which captures sharp dependence on all of $T,K,\lambda$.
@@
For $\beta=1$ (i.e., the slowly-varying setting), the above rate becomes $T^{2/3}\lambda^{1/3}K^{1/3}=T\cdot(\lambda/T)^{1/3}\cdot K^{1/3}$ which is the rate seen in Slivkins (2014) for drift parameter $\delta=\lambda/T$.
@@ 4   Dynamic Regret Upper Bound
As alluded in Subsection 1.2, our main dynamic regret upper bound is achieved by the META algorithm of Suk and Kpotufe (2022) which adapts to so-called significant shifts in best arm. The key idea behind this result is that a significant shift encodes large variation with respect to any $(\beta,\lambda)$-Hölder environment, thus allowing us to recover the rate of Theorem 3. We now recall the notion of a significant shift.
First, we say arm $a$ incurs significant regret [^2] Our definition is slightly different from that of Suk and Kpotufe (2022); all mentioned results hold for either notions. on interval [^3] From here on, we’ll conflate the intervals $[a,b],[a,b)$ for $a,b\in\mathbb{N}$ with the natural numbers contained within. $[s_{1},s_{2}]$ if:
@@
or intuitively if it incurs large dynamic regret. On the other hand, if (1) holds for no interval in a time period, then arm $a$ incurs little regret over that period and is safe to play. Thus, a significant shift is recorded only when there is no safe arm left to play. The following recursive definition captures this.
@@
Let $\tau_{0}=1$. Then, recursively for $i\geq 0$, the $(i+1)$-th significant shift is recorded at time $\tau_{i+1}$, which denotes the earliest time $\tau\in(\tau_{i},T]$ such that for every arm $a\in[K]$, there exists rounds $s_{1}<s_{2},[s_{1},s_{2}]\subseteq[\tau_{i},\tau]$, such that arm $a$ has significant regret (1) on $[s_{1},s_{2}]$.
We will refer to intervals $[\tau_{i},\tau_{i+1}),i\geq 0,$ as significant phases. The unknown number of such phases (by time $T$) is denoted ${\tilde{L}}+1$, whereby $[\tau_{\tilde{L}},\tau_{{\tilde{L}}+1})$, for $\tau_{{\tilde{L}}+1}\doteq T+1,$ denotes the last phase.
Then, a significant shift oracle, which roughly plays arms until they’re unsafe in each significant phase and then restarts at each significant shift, attains a regret bound of (Suk and Kpotufe, 2022, Proposition 1)
@@
The main result of Suk and Kpotufe (2022) is to match the above rate up to log terms without any knowledge of non-stationarity. Their META algorithm estimates when the significant shifts $\tau_{i}$ occur using importance-weighted estimates of the gaps and then restarts an elimination procedure upon detecting an empirical version of a significant shift. The inner workings of the algorithm are beyond the scope of this discussion and surprisingly irrelevant for our result here; the interested reader is deferred to Section 4 of Suk and Kpotufe (2022).
Our main result is that, independent of the algorithm and for any environment, the regret rate Equation 2 inherently captures the minimax rate for smooth non-stationary bandits.
@@
Consider any $(\beta,\lambda)$-Hölder environment over $T$ rounds and let $\{\tau_{i}\}_{i=0}^{{\tilde{L}}}$ be the significant shifts of the environment as in Definition 4. Then, we have:
@@
By Theorem 1 of Suk and Kpotufe (2022), the META algorithm has an expected regret upper bound:
@@
Note that any non-stationary bandit environment over $T$ rounds can be captured by a $(\beta,\lambda)$-Hölder environment for any $\beta>0$ using, e.g., a Lagrange interpolation of the finite data $\{(t/T,\mu_{t}(a))\}_{t\in[T],a\in[K]}$. As $T^{\frac{\beta+1}{2\beta+1}}\cdot\lambda^{\frac{1}{2\beta+1}}\cdot K^{\frac{\beta}{2\beta+1}}\to\sqrt{KT}$ as $\beta\to\infty$, this seems to suggest we can recover a stationary $\sqrt{KT}$ regret rate for any non-stationary environment, which is seemingly a contradiction. However, taking $\beta\to\infty$ will make the bound of Theorem 5 vacuous as there is a constant dependence of $\sqrt{\beta+1}$ on $\beta$. This suggests perhaps the $\sqrt{\beta}$ dependence is unavoidable, and it’s curious if such a dependence can be tightened.

### v2

@@ Abstract
Next, we show this optimal dynamic regret can be attained adaptively, without knowledge of $\beta,\lambda$. To contrast, even with parameter knowledge, upper bounds were only previously known for limited regimes $\beta\leq 1$ and $\beta=2$ (Slivkins, 2014; Krishnamurthy and Gopalan, 2021; Manegueu et al., 2021; Jia et al., 2023). thus, our work resolves open questions raised by disparate threads of the literature.
We also study the problem of attaining faster gap-dependent regret rates in non-stationary bandits. While such rates are long known to be impossible in general (Garivier and Moulines, 2011), we show that environments admitting a safe arm Suk and Kpotufe (2022) allow for much faster rates than the worst-case scaling with $\sqrt{t}$. While previous works in this direction focused on attaining the usual logarithmic regret bounds, as summed over stationary periods, our new gap-dependent rates reveal new optimistic regimes of non-stationarity where even the logarithmic bounds are pessimistic. We show our new gap-dependent rate is tight and that its achievability (i.e., as made possible by a safe arm) has a surprisingly simple and clean characterization within the smooth Hölder class model.
@@ 1   Introduction
In multi-armed bandits (MAB), an agent sequentially chooses actions, from a set of $K$ arms, based on partial and uncertain feedback in the form of (bounded) rewards $Y_{t}(a)$ for past actions $a\in[K]$ (see Bubeck and Cesa-Bianchi (2012); Slivkins (2019); Lattimore and Szepesvári (2020) for general surveys). The goal is to maximize the cumulative reward.
We consider the non-stationary smooth variant of the problem where only mild Hölder class assumptions are made on changes in rewards over time. In fact, this model captures any finite-horizon bandit problem (e.g., via a polynomial interpolation). Additionally, the degree of smoothness (as measured by the Hölder exponent or coefficient of the associated Hölder class) can be considered a more fine-grained measure of non-stationarity in comparison to conventional measures appearing in other works on non-stationary MAB. Indeed, the rates in this model smoothly interpolate between the more parametric $\sqrt{LT}$ rates seen in switching bandits (Garivier and Moulines, 2011), with $L$ switches in rewards over horizon $T$, and the $V^{1/3}T^{2/3}$ rates in terms of the total variation measure $V$ quantifying magnitude of total changes in rewards (Besbes et al., 2019).
@@
Our result is also somewhat surprising since prior approaches (Krishnamurthy and Gopalan, 2021; Manegueu et al., 2021; Jia et al., 2023) all relied on confidence bounds on the magnitude of change in rewards, which arises in estimating the bias in estimating rewards due to non-stationarity. The design of such confidence bounds necessarily requires knowledge of the smoothness. To contrast, the META algorithm randomly schedules periods of fresh exploration, many of which coincide with periods where such bias is minimal, whereas such exploration is deterministically scheduled in the aforementioned prior works, and thus must take into account the bias in estimation. In non-parametric statistics, it’s well known that it’s impossible to design confidence intervals adaptive to unknown smoothness (Low, 1997), ruling out approaches of this kind.
@@ 1.1   Further Discussion on Related Works
Switching bandits was first considered in the adversarial setting by Auer et al. (2002), where a version of EXP3 was shown to attain optimal dynamic regret $\sqrt{LT}$ when tuned with knowledge of the number $L$ of switches. Later works showed similar guarantees in this problem for procedures inspired by stochastic bandit algorithms (Kocsis and Szepesvári, 2006; Yu and Mannor, 2009; Garivier and Moulines, 2011; Mellor and Shapiro, 2013; Liu et al., 2018; Cao et al., 2019). Recently, Auer et al. (2018, 2019); Chen et al. (2019) established the first adaptive and optimal dynamic regret guarantees, without requiring knowledge of $L$. Other non-stationarity measures, such as the aforementioned total variation, or more nuanced counts than $L$ were studied (Suk and Kpotufe, 2022; Abbasi-Yadkori et al., 2023).
@@ 1.2   Contributions
Parameters Studied in Prior WorksAdaptive?Dynamic Regret Upper Bound β(0,1] (Manegueu et al., 2021) NoTβ+12β+1λ12β+1Kβ2β+1
=β1 (Slivkins, 2014; Krishnamurthy and Gopalan, 2021)
 NoT23λ13K13
 β=1,2K=2 (Jia et al., 2023) NoTβ+12β+1λ12β+1
>β0 (this work)
@@
=β1 (Slivkins, 2014; Krishnamurthy and Gopalan, 2021)
@@ 2.1   Preliminaries and Notation
Arm $a$ at round $t$ has random reward $Y_{t}(a)\in[0,1]$ with mean $\mu_{t}(a)$. A (possibly randomized) algorithm $\pi$ selects at each round $t$ some arm $\pi_{t}\in[K]$ and observes reward $Y_{t}(\pi_{t})$. The goal is to minimize the cumulative or total dynamic regret, i.e., the expected regret to the best arm at each round. This is defined as
@@
we will use $R_{\mathcal{E}}(\pi,T)$ to denote the expected regret under an environment $\mathcal{E}$.
@@
Throughout this paper, in theorem statements we will use $C_{0},C_{1},\ldots$ to denote universal constants free of $K,T,\beta,\lambda,\{\mu_{t}(a)\}_{t\in[T],a\in[K]}$. In proofs, universal constants $c_{0},c_{1},\ldots$ will be used.
@@ 2.2   Smooth Non-Stationary Bandits
We say a bandit environment is $(\beta,\lambda)$-Hölder if, for every arm $a\in[K]$, there exists a $(\beta,\lambda)$-Hölder function $f$ such that the gap function (in time) is realized by $f$, i.e. $\delta_{t}(a)=f(t/T)$ for all $t\in[T]$. we will use $\Sigma(\beta,\lambda)$ to denote the class of bandit environments which are $(\beta,\lambda)$-Hölder over $T$ rounds.
@@ 3   Dynamic Regret Lower Bound
We first characterize the minimax regret rate over the class of problems in $\Sigma(\beta,\lambda)$. For comparison, (Jia et al., 2023, Theorem 3.4) already established a lower bound for integer smoothness $\beta\in\mathbb{Z}_{\geq 1}$, $K=2$ arms, and fixed Hölder coefficient $\lambda=1$. Our main novelty here is to show a more comprehensive lower bound which captures sharp dependence on all of $T,K,\lambda$.
@@
For $\beta=1$ (i.e., the slowly-varying setting), the above rate becomes
$T^{2/3}\lambda^{1/3}K^{1/3}=T\cdot(\lambda/T)^{1/3}\cdot K^{1/3},$
which is the rate seen in Slivkins (2014) for drift parameter $\delta=\lambda/T$.
@@ 4   Dynamic Regret Upper Bound
As alluded in Subsection 1.2, our main dynamic regret upper bound is achieved by the META (Meta-Elimination while Ttracking Aarms) algorithm (given here as Algorithm 1 of Suk and Kpotufe (2022) which adapts to so-called significant shifts in best arm. The key idea behind this result is that a significant shift encodes large variation with respect to any $(\beta,\lambda)$-Hölder environment, thus allowing us to recover the rate of Theorem 1. We now recall the notion of a significant shift.
First, we say arm $a$ incurs significant regret [^2] Our definition is slightly different from that of Suk and Kpotufe (2022); all mentioned results hold for either notions. on interval [^3] From here on, we will conflate the intervals $[a,b],[a,b)$ for $a,b\in\mathbb{N}$ with the naturals contained within. $[s_{1},s_{2}]$ if
@@
or intuitively if it incurs large dynamic regret. On the other hand, if (1) holds for no interval in a time window, then arm $a$ incurs little regret over that period and is safe to play. Thus, a significant shift is recorded only when there is no safe arm left to play. Equivalently, this occurs when every arm $a\in[K]$ must satisfy Equation 1 on some interval $[s_{1},s_{2}]$ of arms. Importantly, these notions are independent of the magnitude or smoothness of non-stationarity, or even whether changes in best-arm have occurred. The following recursive definition captures this.
@@
Let $\tau_{0}=1$. Then, recursively for $i\geq 0$, the $(i+1)$-st significant shift is recorded at time $\tau_{i+1}$, which denotes the earliest time $\tau\in(\tau_{i},T]$ such that for every arm $a\in[K]$, there exist rounds $s_{1}<s_{2},[s_{1},s_{2}]\subseteq[\tau_{i},\tau]$, such that arm $a$ incurs significant regret (1) on $[s_{1},s_{2}]$, or else we let $\tau_{i+1}\doteq T+1$ if no such round exists.
We will refer to the intervals $[\tau_{i},\tau_{i+1}),i\geq 0,$ as significant phases. The unknown number of such phases (by time $T$) is denoted ${\tilde{L}}+1$, whereby $[\tau_{\tilde{L}},\tau_{{\tilde{L}}+1})$, for $\tau_{{\tilde{L}}+1}\doteq T+1,$ denotes the last phase.
Figure 1: An example of a non-stationary safe environment where no significant shift occurs because arm $3$ is safe throughout, maintaining small dynamic regret, even while being suboptimal at all times.
Then, a significant shift oracle, which roughly plays arms until they are unsafe in each significant phase and then restarts at each significant shift, attains a regret bound of Suk and Kpotufe (2022, Proposition 1).
@@
The main result of Suk and Kpotufe (2022) is to match the above rate up to log terms without any knowledge of non-stationarity. Their META algorithm estimates when the significant shifts $\tau_{i}$ occur using importance-weighted estimates of the gaps and then restarts an elimination procedure upon detecting an empirical version of a significant shift, based on importance-weighted estimates $\hat{\delta}_{t}(a)$ of the gaps $\delta_{t}(a)$.
The inner workings of the algorithm are beyond the scope of this discussion and surprisingly irrelevant for our result here, as our regret upper bound holds in a blackbox manner for any algorithm acheiving the optimal rate Equation 2 in terms of significant phases. Our main result is that, independent of the algorithm and for any environment, the regret rate Equation 2 inherently captures the minimax rate for smooth non-stationary bandits. Intuitively, this is because a significant phase $[\tau_{i},\tau_{i+1})$ is in fact long enough to bound the number of significant shifts ${\tilde{L}}$ by $T^{\frac{1}{2\beta+1}}\lambda^{\frac{2}{2\beta+1}}K^{-\frac{1}{2\beta+1}}$ which, when plugged into the optimal rate of $\sqrt{{\tilde{L}}KT}$ yields the optimal rate for smooth problems as seen in Theorem 1.
@@
Consider any $(\beta,\lambda)$-Hölder environment over $T$ rounds and let $\{\tau_{i}\}_{i=0}^{{\tilde{L}}}$ be the significant shifts of the environment as in Definition 3. Then, we have
@@
By Theorem 1 of Suk and Kpotufe (2022), the META algorithm has an expected regret upper bound
@@
Note that any non-stationary bandit environment over $T$ rounds can be captured by a $(\beta,\lambda)$-Hölder environment for any $\beta>0$ using, e.g., a Lagrange interpolation of the finite data $\{(t/T,\mu_{t}(a))\}_{t\in[T],a\in[K]}$. As $T^{\frac{\beta+1}{2\beta+1}}\cdot\lambda^{\frac{1}{2\beta+1}}\cdot K^{\frac{\beta}{2\beta+1}}\to\sqrt{KT}$ as $\beta\to\infty$, this seems to suggest we can recover a stationary $\sqrt{KT}$ regret rate for any non-stationary environment, which is seemingly a contradiction. However, taking $\beta\to\infty$ will make the bound of Theorem 2 vacuous as there is a constant dependence of $\sqrt{\beta+1}$ on $\beta$. This suggests perhaps the $\sqrt{\beta}$ dependence is unavoidable, and it is curious if such a dependence can be tightened.
5   More Details About META
1 Input: horizon $T$.
2 Initialize: round count $t\leftarrow 1$.
3 Episode Initialization (setting global variables $\hat{\tau}_{\ell},\operatorname{\mathcal{A}_{{\text{global}}}},B_{s,m}$):
$\hat{\tau}_{\ell}\leftarrow t$.
  // Start of the $\ell$-th episode.
$\operatorname{\mathcal{A}_{{\text{global}}}}\leftarrow[K]$
  // Global candidate arm set.
4 For each $m=2,4,\ldots,2^{\lceil\log(T)\rceil}$ and $s=\hat{\tau}_{\ell}+1,\ldots,T$:
Sample and store $B_{s,m}\sim\text{Bernoulli}\left(\frac{1}{\sqrt{m\cdot(s-\hat{\tau}_{\ell})}}\right)$.
  // Set replay schedule.
[…]
6 if $t<T$ then restart from Line 2 (i.e. start a new episode).
Algorithm 1 Meta-Elimination while Tracking Arms (META)
1 Input: starting round $\operatorname{\it{t}_{{\text{start}}}}$, scheduled duration $m_{0}$.
Initialize: $t\leftarrow\operatorname{\it{t}_{{\text{start}}}}$, $\mathcal{A}_{t}\leftarrow[K]$.
  // $t$ and $\mathcal{A}_{t}$ are global variables.
2 while $t\leq T$ do
       Let $\mathcal{A}_{\text{current}}\leftarrow\mathcal{A}_{t}$
        // Save current candidate arm set.
4       Increment $t\leftarrow t+1$.
5       if $\exists m\text{{\,such that }}B_{t,m}>0$ then
             Let $m\doteq\max\{m\in\{2,4,\ldots,2^{\lceil\log(T)\rceil}\}:B_{t,m}>0\}$.
              // Maximum replay length.
[…]
              // Child replay interrupts parent.
6
7       Evict bad arms:
[…]
10       Restart criterion: if $\operatorname{\mathcal{A}_{{\text{global}}}}=\emptyset$ then RETURN.
11       if $t>\operatorname{\it{t}_{{\text{start}}}}+m_{0}$ then RETURN.
12
Algorithm 2 Base-Alg$(\operatorname{\it{t}_{{\text{start}}}},m_{0})$: Randomized Successive Elimination
We next give a brief, self-contained description of META (Algorithm 1). We first design a base algorithm (Algorithm 2) which works well in safe environments, where there’s no significant change. We’ll then randomly schedule multiple instances of this base algorithm to detect unknown significant shifts.
Base Algorithm: Randomized Successive Elimination.
In such safe environments, there is a safe arm ${a^{\sharp}}$ which does not incur significant regret in the sense of Equation 1. Our base algorithm will be to learn ${a^{\sharp}}$ while eliminating other unsafe arms which satisfy Equation 1. A key idea is that, by definition, the dynamic regret $\sum_{t=1}^{T}\delta_{t}({a^{\sharp}})$ is small $O(\sqrt{T})$ meaning it suffices to minimize the regret to the safe arm $\sum_{t=1}^{T}\delta_{t}({a^{\sharp}},\pi_{t})$. This latter relative regret/gap can be tracked using the importance-weighted estimate
$\hat{\delta}_{t}(a^{\prime},a)\doteq|\mathcal{A}_{t}|\cdot(Y_{t}(a^{\prime})\cdot\boldsymbol{1}\{\pi_{t}=a^{\prime}\}-Y_{t}(a)\cdot\boldsymbol{1}\{\pi_{t}=a\}).$ (3)
Then, at round $t$, we will eliminate arms from an active armset $\mathcal{A}_{t}$ when an empirical analogue of Equation 1 holds using the estimates $\hat{\delta}_{t}(a^{\prime},a)$. In particular, arm $a$ is evicted at round $t$ if, for some fixed $C_{2}>0$, there exists round $t_{0}<t$ such that
$\max_{a^{\prime}\in[K]}\sum_{s=t_{0}}^{t}\hat{\delta}_{s}(a^{\prime},a)>C_{2}\left(\sqrt{K\log(T)\cdot(t-t_{0}+1)}+K\log(T)\right).$ (4)
Notably, in safe environments, the safe arm ${a^{\sharp}}$ is not eliminated as long as estimates are suitably accurate.
Meta-Elimination using Multiple Base Algorithms.
[…]
We play in episodes, with each new episode or restart triggered by the detection of a significant shift. An episode begins by playing according to an ancestor base algorithm scheduled for the remaining rounds. Other descendant base algorithms, called replays, occasionally interrupt the ancestor and become active. Recursively, an active base can further activate its own replays, inducing a hierarchical structure on base algorithms as captured by Figure 2
Figure 2: Shown are two replay durations $m=m_{1}\text{ or }m_{2}$ occurring roughly every ${\sqrt{M/m}}$ rounds following the random schedule of Line 1 of Algorithm 1, where $M$ is the eventual length of an episode. Each replay (blue segment) aims to detect a $1/\sqrt{m}$ magnitude change, i.e., an average dynamic regret $\frac{1}{m}\sum_{t=1}^{m}\delta_{t}(a)$ of order $1/\sqrt{m}$. As a recursive procedure, the replays of Base-Alg  form a parent-child relationship as depicted.
A global arm set $\operatorname{\mathcal{A}_{{\text{global}}}}$ tracks the arms retained by all base algorithms. A restart is triggered when $\operatorname{\mathcal{A}_{{\text{global}}}}$ becomes empty, or when there is no safe arm remaining which aligns with our notion of significant shift (Definition 3).